# Effective 2D Envelope Function Theory for Silicon Quantum Dots

### v1

@@ I Introduction
Silicon spin-based quantum computing is a promising platform for scalable quantum technologies due to its compatibility with existing CMOS infrastructure [1] and the long coherence times of electron spins hosted by quantum dots [3]. The development of scalable silicon quantum devices, however, relies on fast and accurate design–simulation feedback loops [4]. As devices grow in complexity, rapid iteration capabilities over potential geometries become essential. Crucial performance metrics—such as tunnel coupling and, especially, the exchange interaction $J$—are highly sensitive to microscopic details [5] and must be computed numerically with high fidelity. Efficient yet accurate modeling tools [6] are therefore indispensable for guiding experimental design and enabling the scalable control of spin qubits.
@@
Moreover, by including the two lowest eigenstates, which typically correspond to the lowest pair of valley states, we naturally recover also the valley degree of freedom (DOF). This statement is usually justified by the large energy gap to higher orbital and additional valley states, which are typically lifted by strain and interface effects in silicon quantum dots [16]. This valley structure originates from the near-degenerate conduction band minima along the ±$\pm z$ directions in the silicon bandstructure [14] and is essential for modeling realistic quantum dot qubits.
@@ II.1 Modeling Framework
[…]
@@
  • Silicon’s conduction band features six equivalent minima (valleys) located along the 100$\langle$ 100$\rangle$ 100 ⟩ directions in the Brillouin zone. Under confinement and strain, the four in-plane valleys (±$\pm\hat{x},\pm\hat{y}$) are typically raised in energy (see, e.g., [18]), leaving two low-energy valleys along the ±$\pm\hat{z}$ direction. These lead to interference patterns in the wavefunction on the scale of $1/|\mathbf{k}_{0}|$, where ±$\pm\mathbf{k}_{0}=\pm 0.85\times\frac{2\pi}{a_{\text{Si}}}\hat{z}\approx\pm 9.83~\text{nm}^{-1}\hat{z}$ are the valley wavevectors.
@@
$\Psi(\mathbf{r})=\sum_{\mathbf{k}}c_{\mathbf{k}}u_{\mathbf{k}}(\mathbf{r})e^{i\mathbf{k}\cdot\mathbf{r}}\approx F_{+}(\mathbf{r})u_{+\mathbf{k}_{0}}(\mathbf{r})e^{i\mathbf{k}_{0}\cdot\mathbf{r}}+F_{-}(\mathbf{r})u_{-\mathbf{k}_{0}}(\mathbf{r})e^{-i\mathbf{k}_{0}\cdot\mathbf{r}},$ (3)
@@
This approximation assumes that only states near the valley minima contribute. This is equivalent to saying that $F_{\pm}(\mathbf{r})$ does not vary much on the scale of a unit cell. Furthermore,
@@
We now propose an intermediate model between the full Bloch-state approach and the envelope approximation. In this Generalized Envelope Function (GEF) framework, we remove the periodic part $u_{\pm\mathbf{k}_{0}}(\mathbf{r})$ of the Bloch functions while retaining the essential fast oscillations arising from the multi-valley structure. All Bloch-function complexity is abstracted into a single momentum-dependent kernel. We begin by defining the GEF $\psi(\mathbf{r})$ as the spatial representation of the $k$-space coefficients $c_{\mathbf{k}}$. In the discrete form:
@@
where $\Omega$ is the plane wave normalization. In the limit where contributions from only two valleys ±$\pm\mathbf{k}_{0}$ dominate, the function $\psi(\mathbf{r})$ can be approximately decomposed as a superposition of two slowly varying components $F_{\pm}(\mathbf{r})$, each modulated by the respective valley phase factor $e^{\pm i\mathbf{k}_{0}\cdot\mathbf{r}}$. That is,
$\psi(\mathbf{r})\approx F_{+}(\mathbf{r})e^{i\mathbf{k}_{0}\cdot\mathbf{r}}+F_{-}(\mathbf{r})e^{-i\mathbf{k}_{0}\cdot\mathbf{r}}.$ (7)
This highlights that the GEF is a generalization of the traditional envelope function formalism: it retains fast oscillations from valley interference directly in $\psi(\mathbf{r})$, without explicit reference to the periodic part of the Bloch functions. When projected onto a restricted valley subspace, the GEF reduces to the standard envelope function model with valley-dependent envelopes $F_{\pm}$. In the continuum limit, the GEF can be represented as,
@@
Comparing Eq. (6) with the full wavefunction $\Psi(\mathbf{r})=\sum_{\mathbf{k}}u_{\mathbf{k}}(\mathbf{r})e^{i\mathbf{k}\cdot\mathbf{r}}c_{\mathbf{k}}$, we see that the GEF removes only the atomically varying Bloch functions $u_{\mathbf{k}}(\mathbf{r})$, but still retains the oscillatory features at $\mathbf{k}_{0}$. Thus, $\psi(\mathbf{r})$ is a smooth but valley-aware approximation to the full wavefunction. The governing equation for the GEF is an effective Schrödinger equation,
@@
As a model that should capture all relevant features of the bandstructure, we choose to approximate $\epsilon(\mathbf{k})$ as a sum of two anisotropic Gaussians centered at the valley minima near ±$\pm k_{0}\hat{z}$:
@@
For non-valley-resolved calculations, we only consider a single minimum of the bandstructure and may therefore resort to a simpler representation of the kinetic energy operator, valid within the effective mass approximation. In this case, the Hamiltonian has the standard form
^$\hat{\mathcal{H}}=-\frac{\hbar^{2}}{2m_{t}}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)-\frac{\hbar^{2}}{2m_{z}}\frac{\partial^{2}}{\partial z^{2}}.$ (12)
This form is justified as here, we approximate the bandstructure at the valley minimum harmonically and can therefore identify the kinetic term with differential operators.
@@ II.2 Born-Oppenheimer-Inspired Dimensional Reduction
The validity of these approximations will be addressed in later sections.
@@
[…]
@@ II.3 Effective 2D Hamiltonian in Charge Space
[…]
@@ II.4 Effective Hamiltonian in Charge and Valley Space — 2D Multivalley Envelope Function Theory
In this subsection, we evaluate the projected Hamiltonian (17) using the two lowest vertical eigenstates, which—as discussed earlier—form a near-degenerate valley pair centered at ±$\pm k_{0}\hat{z}$. To be permitted to use the general form of Eq. (17), we must first assess whether the Hamiltonian is approximately separable as assumed in Eq. (13) and whether the form of the kinetic energy operator is in accordance with Eq. (14). Starting from the full Hamiltonian (17), and using the anisotropic bandstructure model of Eq. (11), we proceed by making a further simplification: we approximate the in-plane ($x$-$y$) bandstructure quadratically at low energies, i.e., we assume that the bandstructure can be written as a parabolic function in $x$ and $y$, superposed with the double well function in $z$.
Under these assumptions, the Hamiltonian is approximately separable, and we are justified in applying the dimensional reduction framework from the previous section. A complication arises, however, from the fact that the valley structure of the vertical eigenstates is highly sensitive to interface disorder, especially in realistic Si/SiGe or Si/SiO2 devices. To avoid spatial fluctuations in the basis, and to suppress the coupling terms in Eq. (18), we choose to work in a basis where we fix the valley-dependent terms in the vertical eigenfunctions (but still allow the vertical envelope function to vary as a function of $x$ and $y$); we call this the ‘fixed valley-spinor basis’ $\{\tilde{\chi}_{+},\tilde{\chi}_{-}\}$ (see Appendix B for details of how it is constructed). In this basis, the coupling terms (18) reflect only the $xy$-variation of the envelope function; provided this can be neglected, the effective 2D multivalley Hamiltonian becomes
@@
where $\epsilon_{g}(x,y)$ is the local valley ground state energy and $|\Delta(x,y)|$ is the position-dependent valley splitting. Finally, $\phi(x,y)=\arg\Delta(x,y)$ denotes the local valley phase, defined through the complex phase of the intervalley coupling. The matrix in the second term of Eq. (23) captures the essential valley physics emerging from interface roughness and confinement-induced mixing. Both Eq. (23) and the form of $\hat{H}_{z}^{(+,-)}(x,y)$ are derived in Appendix B in greater detail. Appendix B also contains a prescription on how to determine the quantities $\phi(x,y)$ and $\Delta(x,y)$.
@@ III.2 Solutions to the stationary Schrödinger Equation
The kinetic energy operator is applied using the anisotropic band dispersion relation $\epsilon(\mathbf{k})$, defined in Eq. (11), via
$\hat{T}=\mathcal{F}^{-1}\left[\epsilon(\mathbf{k})\mathcal{F}[\psi]\right],$
@@ III.3 Valley Coupling Kernel Approximation
To represent inter-valley coupling in a numerically stable and physically grounded way, we model $C_{0}(\mathbf{q})$ using negative Gaussian functions centered at the critical coupling points:
$C_{0}(\mathbf{q})=1+A_{G}\exp\left(-\frac{|\mathbf{q}-2\mathbf{k}_{0}|^{2}}{2\sigma^{2}}\right)+A_{G}\exp\left(-\frac{|\mathbf{q}+2\mathbf{k}_{0}|^{2}}{2\sigma^{2}}\right),$
where $\sigma=2.0\,\mathrm{nm}^{-1}$ defines the Gaussian width and $A_{G}=-1.26$ sets the amplitude. This form ensures the kernel is approximately unity away from the coupling points and strongly suppressed near them:
$C_{0}(\pm 2\mathbf{k}_{0})=1+A_{G}=-0.26.$
The Gaussian suppression provides a smooth, physically motivated modulation in momentum space, avoiding the numerical artifacts associated with step-function models and enabling efficient evaluation via FFTs. For the $z$-valleys, the coupling points are located at:
$\mathbf{q}_{\pm}=[0,0,\pm 2k_{0,z}]\approx[0,0,\pm 19.58]\,\mathrm{nm}^{-1}.$
This Gaussian approximation captures both the locality and strength of inter-valley interactions, and typically affects only a small fraction of the momentum grid (2–5%), consistent with the narrow momentum-space character of valley coupling.
@@ III.4 Computational Cost Analysis
2D BO (1e) 60×1260$\times$ 1260 × 12 720720720 0.010.010.01 111
3D Full (1e) 60×12×2460$\times$ 12$\times$ 2460 × 12 × 24 17,28017,28017 , 280 0.260.260.26 $10^{2}$
2D BO (2e) $(60\times 12)^{2}$ 518,400518{,}400518 , 400 7.917.917.91 $10^{3}$
3D Full (2e) $(60\times 12\times 24)^{2}$ 298,598,400298{,}598{,}400298 , 598 , 400 455645564556 $10^{7}$
@@
[…]
2D BO Valley (2e) $(60\times 12\times 2)^{2}$ 8,294,4008{,}294{,}4008 , 294 , 400 127127127 $10^{3}$
3D Full Valley (2e) $(60\times 12\times 100)^{2}$ $5.18\times 10^{9}$ 791027910279102 infeasible
@@ IV.1.1 Device Model
We model a double quantum dot (DQD) device consisting of metallic gates positioned above a silicon heterostructure. The device geometry is defined within a computational domain of 160×160×50160$\times$ 160$\times$ 50160 × 160 × 50 nm3 (see Figure 4(b)). The gate architecture comprises five distinct electrodes: three plunger gates and two confinement gates, as illustrated in Fig. 1. The left and right plunger gates (P1 and P3) are positioned at $x=-30$ nm and $x=30$ nm, respectively, each with dimensions of 25×60×1025$\times$ 60$\times$ 1025 × 60 × 10 nm3. The central barrier gate (P2) is located at $x=0$ nm with dimensions 25×60×1025$\times$ 60$\times$ 1025 × 60 × 10 nm3 and is positioned 5 nm closer to the substrate surface than the other two gates to enhance electrostatic control over the interdot tunnel coupling. Two confining gates (C1 and C2) provide lateral confinement, positioned at $y=\pm 50$ nm with extended dimensions of 160×30×10160$\times$ 30$\times$ 10160 × 30 × 10 nm3. To obtain the potential for a given gate voltage arrangement, we solve the Laplace equation as outlined in Sec. III. We solve the latter on a grid with a resolution of 1 nm in the $x$ and $y$ directions and 0.5 nm in the $z$ direction.
For charge-only calculations, we investigate two device configurations: a flat interface structure and a stepped heterostructure. The stepped configuration features three distinct interface sections with heights varying between 0.5-0.5- 0.5 nm and +0.5+0.5+ 0.5 nm, representing realistic fabrication variations that arise during epitaxial growth or etching processes. In particular, variations at this magnitude should be expected as these have been observed in tunnling electron microscopy measurements of real devices, see [27]. For this section, we assume that the material stack consists of SiO2 ($\epsilon_{r}=3.9$) above the interface and a silicon ($\epsilon_{r}=11.9$) substrate. These varying permittivities are considered in the solution of the Laplace equation. Additionally, there is a conduction band offset of 3 eV between the Si and SiO2 regions. We add this value to the gate-defined electrostatic potential in the SiO2 domain (all space above $z>0$) to confine the quantum dot in the $z$-direction in a realistic manner.
@@
Figure 2: Cross-sectional view of the stepped heterostructure double quantum dot along the $xz$-plane at $y=0$. The stepped interface lies beneath the three main gates (left plunger, barrier, and right plunger). It consists of three sections: Section 1 ($x\in[-80,-10]$ nm) at $z=0.5$ nm, Section 2 ($x\in[-10,+10]$ nm) at $z=-0.5$ nm, and Section 3 ($x\in[+10,+80]$ nm) at $z=0$ nm. Silicon lies below the interface while SiO2 is above.

### v2

@@ I Introduction
Silicon spin-based quantum computing is a promising platform for scalable quantum technologies due to its compatibility with existing CMOS infrastructure [1] and the long coherence times of electron spins hosted by quantum dots [2]. The development of scalable silicon quantum devices, however, relies on fast and accurate design–simulation feedback loops [3]. As devices grow in complexity, rapid iteration capabilities over potential geometries become essential. Crucial performance metrics—such as tunnel coupling and, especially, the exchange interaction JJ—are highly sensitive to microscopic details [4] and must be computed numerically with high fidelity. Efficient yet accurate modeling tools [5] are therefore indispensable for guiding experimental design and enabling the scalable control of spin qubits.
@@
Moreover, by including the two lowest eigenstates, which typically correspond to the lowest pair of valley states, we naturally recover also the valley degree of freedom (DOF). This statement is usually justified by the large energy gap to higher orbital and additional valley states, which are typically lifted by strain and interface effects in silicon quantum dots [15]. This valley structure originates from the near-degenerate conduction band minima along the $\pm z$ directions in the silicon bandstructure [13] and is essential for modeling realistic quantum dot qubits.
@@ II.1 Modeling Framework
We begin with a brief review of how to describe the quantum dot wavefunction in a gate-defined silicon quantum dot, along with existing methods to simplify its representation. In general, the wavefunction of an electron in a silicon heterostructure experiences a highly non-trivial potential consisting of three components: the silicon crystal potential, an abrupt dielectric interface between silicon and the barrier material (e.g., silicon-oxide), and the external electrostatic potential generated by gate electrodes. Hence, the total Hamiltonian describing the system in 3DD can be written as:
@@
  • Silicon’s conduction band features six equivalent minima (valleys) located along the $\langle 100\rangle$ directions in the Brillouin zone. Under confinement and strain, the four in-plane valleys ($\pm\hat{x},\pm\hat{y}$) are typically raised in energy (see, e.g., [17]), leaving two low-energy valleys along the $\pm\hat{z}$ direction. These lead to interference patterns in the wavefunction on the scale of $1/|\mathbf{k}_{0}|$, where $\pm\mathbf{k}_{0}=\pm 0.85\times\frac{2\pi}{a_{\text{Si}}}\hat{z}\approx\pm 9.83~\text{nm}^{-1}\hat{z}$ are the valley wavevectors.
@@
$\Psi(\mathbf{r})=\sum_{\mathbf{k}}u_{\mathbf{k}}(\mathbf{r})e^{i\mathbf{k}\cdot\mathbf{r}}c_{\mathbf{k}}\approx F_{+}(\mathbf{r})u_{+\mathbf{k}_{0}}(\mathbf{r})e^{i\mathbf{k}_{0}\cdot\mathbf{r}}+F_{-}(\mathbf{r})u_{-\mathbf{k}_{0}}(\mathbf{r})e^{-i\mathbf{k}_{0}\cdot\mathbf{r}},$ (3)
@@
This approximation assumes that only states near the valley minima contribute. This is equivalent to saying that $F_{\pm}(\mathbf{r})$ does not vary much on the scale of a unit cell. Furthermore, in (3) we used,
@@
We now propose an intermediate model between the full Bloch-state approach and the envelope approximation. In this Generalized Envelope Function (GEF) framework, we remove the periodic part $u_{\pm\mathbf{k}_{0}}(\mathbf{r})$ of the Bloch functions while retaining the essential fast oscillations arising from the multi-valley structure. All Bloch-function complexity is abstracted into a single momentum-dependent kernel. We define the GEF as the spatial representation of the $k$-space coefficients. In a finite normalization volume $\Omega$,
@@
and, in the continuum limit,
@@
Comparing with the full wavefunction $\Psi(\mathbf{r})=\sum_{\mathbf{k}}u_{\mathbf{k}}(\mathbf{r})\,e^{i\mathbf{k}\cdot\mathbf{r}}\,c_{\mathbf{k}},$ the GEF removes only the atomically varying factors $u_{\mathbf{k}}(\mathbf{r})$ while keeping the valley-phase oscillations carried by $e^{i\mathbf{k}\cdot\mathbf{r}}$.
When two valleys at $\pm\mathbf{k}_{0}$ dominate, $\psi$ admits the standard two-valley decomposition with slowly varying envelopes,
$\psi(\mathbf{r})\;\approx\;F_{+}(\mathbf{r})\,e^{+i\mathbf{k}_{0}\cdot\mathbf{r}}\;+\;F_{-}(\mathbf{r})\,e^{-i\mathbf{k}_{0}\cdot\mathbf{r}},$ (8)
so that, upon restriction to the valley subspace, the GEF reduces to the familiar envelope-function model with valley-dependent envelopes $F_{\pm}(\mathbf{r})$. The governing equation for the GEF is an effective Schrödinger equation,
@@
As a model that should capture all relevant features of the bandstructure, we choose to approximate $\epsilon(\mathbf{k})$ as a sum of two anisotropic Gaussians centered at the valley minima near $\pm k_{0}\hat{z}$:
@@
For non-valley-resolved calculations, we only consider a single minimum of the bandstructure and may therefore resort to a simpler representation of the kinetic energy operator, valid within the effective mass approximation. In this case, the kinetic energy operator exhibits the standard form
$\hat{\mathcal{T}}=-\frac{\hbar^{2}}{2m_{t}}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)-\frac{\hbar^{2}}{2m_{z}}\frac{\partial^{2}}{\partial z^{2}},$ (12)
which is justified as here, we approximate the conduction band dispersion at the valley minimum harmonically and can therefore identify the kinetic term with differential operators.
@@ II.2 Born-Oppenheimer-Inspired Dimensional Reduction
The validity of these approximations in our cases of interest will be addressed in later sections.
@@
$D^{(0)}_{m^{\prime}m}(x,y)$ $=\int dz\,\chi^{*}_{m^{\prime}}\,(\partial_{x}^{2}+\partial_{y}^{2})\chi_{m},$ (18)
$D^{(1,x)}_{m^{\prime}m}(x,y)$ $=2\int dz\,\chi^{*}_{m^{\prime}}\,\partial_{x}\chi_{m},$
$D^{(1,y)}_{m^{\prime}m}(x,y)$ $=2\int dz\,\chi^{*}_{m^{\prime}}\,\partial_{y}\chi_{m}.$
@@ II.3 Effective 2D Hamiltonian in Charge Space
where the new effective 2DD potential, $V_{\text{eff}}(x,y)$, is simply the sum of the in-plane potential $V_{xy}(x,y)$ and the vertical ground-state energy $\epsilon_{0}(x,y)$. Equation (21) defines a compact and computationally efficient 2D model for describing charge dynamics in quantum dots, with valley effects neglected by construction.
@@ II.4 Effective Hamiltonian in Charge and Valley Space — 2D Multivalley Envelope Function Theory
In this subsection, we evaluate the projected Hamiltonian (17) using the two lowest vertical eigenstates, which—as discussed earlier— usually form a near-degenerate valley pair. To be permitted to use the general form of Eq. (17), we must first assess whether the Hamiltonian is approximately separable as assumed in Eq. (13) and whether the form of the kinetic energy operator is in accordance with Eq. (14). Starting from the full Hamiltonian (17), and using the anisotropic bandstructure model of Eq. (11), we proceed by making a further simplification: we approximate the in-plane ($x$-$y$) bandstructure quadratically at low energies, i.e., we assume that the bandstructure can be written as a parabolic function in $x$ and $y$, superposed with the double well function in $z$. Under these assumptions, the Hamiltonian is approximately separable, and we are justified in applying the dimensional reduction framework discussed in the previous section.
A complication, however, arises from the fact that the valley structure of the vertical eigenstates is highly sensitive to interface disorder, especially in realistic Si/SiGe or Si/SiO2 devices. To avoid spatial fluctuations in the basis, and thereby suppressing the coupling terms in Eq. (18), we choose to work in a basis where we fix the valley-dependent terms in the vertical eigenfunctions (but still allow the vertical envelope function to vary as a function of $x$ and $y$); we call this the ‘fixed valley-spinor basis’ $\{\tilde{\chi}_{+},\tilde{\chi}_{-}\}$ (see Appendix B for details of how it is constructed). As this basis is different from the eigenbasis of the local vertical Hamiltonian $\hat{H}_{z}$ we are required to express the diagonal $\epsilon_{m}(x,y)\delta_{m^{\prime}m}$ term in (17) in it. We call this rotated representation of $\epsilon_{m}(x,y)\delta_{m^{\prime}m}$, $\hat{H}_{z}^{(+,-)}(x,y)_{\nu^{\prime}\nu}$. In this fixed valley basis, the coupling terms (18) manifest only in the $xy$-variation of the envelope function. Provided that coupling to higher energy states can be neglected, the effective 2D multivalley Hamiltonian becomes
@@
where $\epsilon_{g}(x,y)$ is the local valley ground state energy and $|\Delta(x,y)|$ is the position-dependent valley splitting. Finally, $\phi(x,y)=\arg\Delta(x,y)$ denotes the local valley phase, defined through the complex phase of the intervalley coupling. The matrix in the second term of Eq. (23) captures the essential valley physics emerging from interface roughness and confinement-induced mixing. Both Eq. (23) and the form of $\hat{H}_{z}^{(+,-)}(x,y)$ are derived in Appendix B in greater detail. Finally, Appendix C contains a prescription on how to determine the quantities $\phi(x,y)$ and $\Delta(x,y)$.
@@ III.2 Solutions to the stationary Schrödinger Equation
The kinetic energy operator, defined from the band dispersion in Eq. (11), acts in real space via the Fourier operators as
$\hat{T}\,\psi\;=\;\big(\mathcal{F}^{-1}\,\epsilon\,\mathcal{F}\,\psi\big)(\mathbf{r}).$ (26)
@@ III.3 Valley Coupling Kernel Approximation
From earlier works (see Ref. [22]) we can extract that the intervalley coupling value is approximately 0.26-0.26. Since normalization requires $C_{0}(\mathbf{0})=1$, we interpolate between these values by modeling $C_{0}(\mathbf{q})$ as the sum of two negative Gaussians centered at the critical coupling points $\pm 2\mathbf{k}_{0}$:
$C_{0}(\mathbf{q})=1+A_{G}\exp\!\left(-\frac{|\mathbf{q}-2\mathbf{k}_{0}|^{2}}{2\sigma^{2}}\right)+A_{G}\exp\!\left(-\frac{|\mathbf{q}+2\mathbf{k}_{0}|^{2}}{2\sigma^{2}}\right),$ (27)
with Gaussian width $\sigma=2.0~\mathrm{nm}^{-1}$ and amplitude $A_{G}=-1.26$. This choice ensures
$C_{0}(\pm 2\mathbf{k}_{0})\simeq 1+A_{G}=-0.26,$ (28)
because the cross-Gaussian contribution at $\mathbf{q}=\pm 2\mathbf{k}_{0}$ is exponentially suppressed by $\exp\!\big[-(4k_{0})^{2}/(2\sigma^{2})\big]\ll 1$ for the parameters used, while
$C_{0}(\mathbf{0})=1+2A_{G}\exp\!\big[-(2k_{0})^{2}/(2\sigma^{2})\big]\approx 1$ (29)
due to the negligible Gaussian tails at the origin. The resulting kernel is therefore close to unity away from the coupling points and suppressed near $\pm 2\mathbf{k}_{0}$, providing a smooth, physically motivated modulation in momentum space that avoids the Gibbs-type artifacts of step-function models and remains efficient to evaluate.
@@ III.4 Computational Cost Analysis
2D BO (1e) $60\times 12$ $720$ $0.01$ 11
3D Full (1e) $60\times 12\times 24$ $17,280$ $0.26$ $10^{2}$
2D BO (2e) $(60\times 12)^{2}$ 518,400518{,}400 $7.91$ $10^{3}$
3D Full (2e) $(60\times 12\times 24)^{2}$ 298,598,400298{,}598{,}400 $4556$ $10^{7}$
@@
2D BO Valley (1e) $60\times 12\times 2$ 1,4401{,}440 $0.04$ 11
3D Full Valley (1e) $60\times 12\times 100$ 72,00072{,}000 $1.10$ $10^{2}$
2D BO Valley (2e) $(60\times 12\times 2)^{2}$ 8,294,4008{,}294{,}400 $127$ $10^{3}$
3D Full Valley (2e) $(60\times 12\times 100)^{2}$ $5.18\times 10^{9}$ $79102$ infeasible
@@ IV.1.1 Device Model
We model a double quantum dot (DQD) device consisting of metallic gates positioned above a silicon heterostructure. The device geometry is defined within a computational domain of $160\times 160\times 50$ nm3 (see Figure 4(b)). The gate architecture comprises five distinct electrodes: three plunger gates and two confinement gates, as illustrated in Fig. 1. The left and right plunger gates (P1 and P3) are positioned at x=30x=-30 nm and $x=30$ nm, respectively, each with dimensions of $25\times 60\times 10$ nm3. The central barrier gate (P2) is located at $x=0$ nm with dimensions $25\times 60\times 10$ nm3 and is positioned 5 nm closer to the substrate surface than the other two gates to enhance electrostatic control over the interdot tunnel coupling. Two confining gates (C1 and C2) provide lateral confinement, positioned at $y=\pm 50$ nm with extended dimensions of $160\times 30\times 10$ nm3. To obtain the potential for a given gate voltage arrangement, we solve the Laplace equation as outlined in Sec. III. We solve the latter on a grid with a resolution of 1 nm in the $x$ and $y$ directions and 0.5 nm in the $z$ direction.
For charge-only calculations, we investigate two device configurations: a flat interface structure and a stepped heterostructure. The stepped configuration features three distinct interface sections with heights varying between 0.5-0.5 nm and +0.5+0.5 nm, representing realistic fabrication variations that arise during epitaxial growth or etching processes. In particular, variations at this magnitude should be expected as these have been observed in tunnling electron microscopy measurements of real devices, see [26]. For this section, we assume that the material stack consists of SiO2 ($\epsilon_{r}=3.9$) above the interface and a silicon ($\epsilon_{r}=11.9$) substrate. These varying permittivities are considered in the solution of the Laplace equation. Additionally, there is a conduction band offset of 3 eV between the Si and SiO2 regions. We add this value to the gate-defined electrostatic potential in the SiO2 domain (all space above $z>0$) to confine the quantum dot in the $z$-direction in a realistic manner.
@@
Figure 2: Cross-sectional view of the stepped heterostructure double quantum dot along the $xz$-plane at $y=0$. The stepped interface lies beneath the three main gates (left plunger, barrier, and right plunger). It consists of three sections: Section 1 ($x\in[-80,-10]$ nm) at $z=0.5$ nm, Section 2 ($x\in[-10,+10]$ nm) at z=0.5z=-0.5 nm, and Section 3 ($x\in[+10,+80]$ nm) at $z=0$ nm. Silicon lies below the interface while SiO2 is above.